If A + B =90 0 , then find the value of sin 2 A + sin 2 B.

If A + B =900, then find the value of sin2 A + sin2 B.
 

Grade:10

1 Answers

Navdeep Singh
8 years ago
sin2A+sin2B=sin2A+1-cos2B
                  =1-(cos2B-sin2A)
                  =1-(cos(A+B)cos(A-B))
                   =1-(cos(900)cos(A-B))                            ;(A+B=900)
                    =1-(0)                                                          ;cos(900)=0
                    =1